OpenStudy (anonymous):

The given is: angle 1 is an exterior angle of triangle ABC Prove: Measure of angle 1=m

OpenStudy (mathstudent55):



OpenStudy (mathstudent55):

Statement 1: (copy given) Reason: given Statement 2: m<1 + m<2 = 180 Reason: def of supplementary angles Statement 3: m<A + m<B + m<2 = 180 Reason: the sum of the measures of the angles of a triangle equals 180 Statement 4: m<1 + m<2 = m<A + m<B + m<2 Reason: substitution Statement 5: m<1 = m<A + m<B Reason: subtraction property of equality